# A Fast, Analytic Empirical Model of the Gaia Data Release 3 Astrometric Orbit Catalog Selection Function

### v1

@@ 4 Empirical modeling of Gaia NSS model cascade
With a binary population model in hand, the next step is to reproduce the modeling cascade described in §2. In this section, we describe how we estimate and implement the filters $\texttt{ruwe}>1.4$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, and $a_{0}/\sigma_{a_{0}}>\max(5,\sqrt{P_{orb}/\textrm{days}})$. In addition, we also need a way to account for the fact that 1) Gaia DR3 is based on only 34 months of data, 2) orbital solutions are “lost” higher up in the model cascade when they are fit with an acceleration solution, and 3) orbital periods of 1 year are very difficult to distinguish from parallax. We describe the empirical filters we apply to account for these effects.
4.1 Empirical ruwe  model
@@ 4.1.1 Analytic ruwe  prediction
$\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1},$ (3)
where $\sigma_{AL}(G)$ is the along-scan centroiding error, which is a function of the source magnitude $G$.
Here, as well as in the remainder of this paper, we adopt the per-CCD centroiding error from Figure 3 of Holl et al. (2023). As described in Lindegren (2018), for $12\lesssim G\lesssim 17$ the formal centroiding precision is limited by photon noise (leading to Poisson errors), for $G\gtrsim 17$ background noise dominates and the error rises more quickly than Poisson, and for $G\lesssim 12$ the CCD gates limit the number of photons and the centroiding precision is independent of magnitude. The $\sigma_{AL}(G)$ we adopt from Holl et al. (2023) is empirically determined from Gaia DR3 data, and is significantly larger ($>0.3$ dex) than the formal error for bright ($G\gtrsim 12)$ sources.
@@
$\delta\theta=a_{0}\sqrt{1-\frac{\sin^{2}i}{2}-\frac{3+\sin^{2}i(\cos^{2}\phi_{\nu}-2)}{4}e^{2}},$ (4)
where $\phi_{\nu}$ is the azimuthal viewing angle of the binary. Note that the derivation of Equation 4 assumes that there are many observations sampling the full orbital period of the binary during the survey, such that the observations can be time-averaged over the orbit.
For an isotropic distribution of $\phi_{\nu}$, the average value of $\cos^{2}\phi_{\nu}$ is 0.5. Fixing $\cos^{2}\phi_{\nu}$ to 0.5, the astrometric excess noise can be re-written
@@
Figure 2: Using the NSS astrometric orbits catalog, we validate the scaling relations between the approximation of the astrometric excess noise $\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1}$ and the semimajor axis of the photocenter orbit $a_{0}$, eccentricity $e$, and orbital inclination $i$. The left column shows the approximation of $\delta\theta$ vs. $a_{0}$, with the color of the point denoting $|\cos i|$ (top panel) or eccentricity of the orbit (bottom panel). There is a linear trend between $a_{0}$ and $\delta\theta$ where the slope is less than unity, as expected. In addition, there is a clear gradient in the colors of the points: averaging over the orbit, face-on and circular orbits $(|\cos i|\approx 1,e\approx 0)$ have larger astrometric excess noise at fixed $a_{0}$ than edge-on and highly eccentric orbits $(|\cos i|\approx 0,e\approx 1)$. The right column is the same as the left column, but includes the dependence of $e$ and $i$ on $\delta\theta$, i.e. shows the approximation of $\delta\theta$ vs. $a_{0}\sqrt{(1+\cos^{2}i)(1-3e^{2}/4)}$. The trends in the colors of the points mostly disappear (the $|\cos i|$ and $e$ dependence), validating the scaling relation. There are still some minor trends visible in $|\cos i|$, but Gaia Collaboration et al. (2023b) showed that the sensitivity of circular orbits deviates slightly from the the expected $1+\cos^{2}i$, so this is not surprising. Note that there are very few red points in the bottom panels because there are very few high-eccentricity orbits.
@@ 4.1.2 Empirical correction to analytic ruwe  prediction
The predicted astrometric excess noise in Equation 5 and ruwe  in Equation 6 will generally be overestimates of the astrometric excess noise and ruwe  reported by Gaia. The derivation of Equation 5 assumes that the binary signal has not been erroneously absorbed into other astrometric terms (i.e., position, parallax, proper motion, or satellite pointing noise). Penoyre et al. (2020) predicted that this erroneous absorption would happen frequently, often lowering the reported astrometric excess noise and ruwe. With the NSS catalog, we confirm that the predicted astrometric excess noise $\delta\theta$ tends to be smaller than the actual $\delta\theta$ in the catalog by $\approx 60-70\%$ (Figure 3).
We calculate an empirical correction to rescale the analytic ruwe  prediction to match the actual ruwe  using the NSS catalog. The number of visibility periods used in the solution ($n_{vis}$) affects the size of the erroneous absorption. For smaller $n_{vis}$, the average size of the erroneous absorption is larger, and there is more scatter about this average; for larger $n_{vis}$, the average size of the erroneous absorption is smaller, and there is less scatter about this average (Figure 3).
To calculate the correction factor, we split the NSS astrometric orbit solutions into three bins based on the number of visibility periods used: $n_{vis}<23$, $2523\leq n_{vis}<$ 25, and $n_{vis}\geq 25$. These bins were chosen since $n_{vis}$ is a bimodal distribution for the NSS astrometric orbits catalog, with peaks at $n_{vis}=18$ and 27. In each of these bins, a Gaussian is fit to the distribution of the ratio of the actual $\delta\theta$ to the analytically predicted $\delta\theta$ from Equation 5. The mean $\mu$ and standard deviation $\sigma$ for each of the bins are $\mu=0.574$, $\sigma=0.133$ for $n_{vis}<23$; $\mu=0.606$, $\sigma=0.116$ for $2523\leq n_{vis}<$ 25, and $\mu=0.643$, $\sigma=0.099$ for $n_{vis}\geq 25$. To apply this correction, a sample is drawn from the appropriate Gaussian and used to rescale the analytic estimate $\delta\theta$, and in turn ruwe, since $\texttt{ruwe}=\sqrt{1+(\delta\theta/\sigma_{AL})^{2}}$ (Equations 4 and 5). We note that we chose not to directly calculate the correction using the ruwe  reported in the NSS astrometric catalog, because the hard limit $\texttt{ruwe}>1.4$ makes the distribution non-Gaussian and would be more difficult to model than using $\delta\theta$.
Figure 3: Comparison of the astrometric excess noise $\delta\theta$ for sources in the NSS astrometric orbit catalog, vs. the analytic predicted $\delta\theta$ of those sources calculated using the photocenter orbit semimajor axis, eccentricity, and inclination from the NSS catalog and Equation 5. We perform the comparison by the number of visibility periods used $n_{vis}$ in the solution. The top panels show scatter plots of the actual $\delta\theta$ vs. analytic predicted $\delta\theta$, with the 1:1 relation is denoted by the orange line. The bottom panels show the ratio of the actual to analytic predicted $\delta\theta$, with the best-fit Gaussian shown in the black line. The predicted astrometric excess noise is systematically higher than the actual value reported in the NSS catalog. The actual values reported are typically only $\approx 60-70$% of the predicted values and do not depend on the size of $\delta\theta$; the values are more severely underestimated when smaller numbers of visibility periods are used, and the scatter about the average is also larger. Some of the astrometric excess noise being erroneously absorbed into other parts of the astrometric model is likely one contributing factor to the underestimation of $\delta\theta$.
4.2 Empirical parallax and photocenter orbit semimajor axis error
We construct an empirical error model for the parallax and photocenter orbit semimajor axis using the NSS catalog.
The parallax error $\sigma_{\varpi}$ and photocenter orbit semimajor axis $\sigma_{a_{0}}$ are expected to be related to the centroiding error in the along-scan direction, $\sigma_{AL}$, and the number of visibility periods used, $N_{vis}$, according to
$\sigma_{\varpi},\sigma_{a_{0}}\propto\frac{\sigma_{AL}(G)}{\sqrt{N_{vis}}}.$ (7)
Using the number of visibility periods provides a more reliable estimate for position and proper motion uncertainties than the total number of good along-scan observations (Lindegren et al., 2018).
To obtain an estimate for $N_{vis}$ in the population synthesis, we use the Gaia Observation Forecast Tool (GOST) [^7] https://gaia.esac.esa.int/gost/, which predicts when any given part of the sky crosses Gaia’s focal plane. To calculate $N_{vis}$, we evaluate the number of predicted scans at 49152 evenly spaced points on the sky, discard a random 10% of the predictions to mimic losses due to dead time [^8] The GOST page suggests 20% would be more appropriate, but we find 10% gives good agreement with observations., then calculate how many of the remaining predicted observations are separated by at least 4 days.
We first attempted to find the constant of proportionality in Equation 7 relating $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ to $\sigma_{AL}(G)/\sqrt{N_{vis}}$. However, we find there is likely an additional unaccounted-for variable in the magnitude of the source (Appendix A).
Thus, we instead rely on the uncertainties from the NSS catalog to make the estimates of $\sigma_{\varpi}$ and $\sigma_{a_{0}}$. In bins of 0.2 mag and integer values of visibility periods, we calculate the median and standard deviation $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ in the NSS catalog. As expected, having more visibility periods and/or having a source magnitude $8\lesssim G\lesssim 13$ (where $\sigma_{AL}$ is smallest) tends to result in smaller values of $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ (Figure 4). We find the standard deviation is roughly constant when the bin is well-sampled; we thus adopt a constant standard deviation of 0.12 mas for $\log_{10}\sigma_{\varpi}$ and 0.37 mas for $\log_{10}\sigma_{a_{0}}$. We also calculate the minimum value of $\log_{10}\sigma_{\varpi}$ and $\log_{10}\sigma_{a_{0}}$ in each bin.
Figure 4: Median parallax error (left) and photocenter orbit semimajor axis error (right) interpolated using the NSS catalog. The black dots show the location of the grid points in NSS; the inferred interpolated values are shown as the colored boxes. Note the color bar scale is different in the two panels.
The median, standard deviation, and minimum values are used as the mean $\mu$, standard deviation $\sigma$, and minimum $a$, respectively, in a truncated normal distribution
$f(x;\mu,\sigma,a,b)=\begin{dcases}\frac{1}{\sigma}\frac{\phi(x;\mu,\sigma)}{\Phi(a;\mu,\sigma)-\Phi(b;\mu,\sigma)},&a\leq x\leq b\\ 0,&\text{otherwise},\end{dcases}$ (8)
where $\phi$ and $\Phi$ are the distributions of a normal PDF and a normal CDF, respectively. The truncation on the low end prevents artificially low errors; we do not apply the truncation on the higher end (hence the parameter $b$ in Equation 8 is infinite; in practice, it is set as a large number like 999).
We then interpolate across this grid of values to determine the median $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ for a given magnitude and number of visibility periods used. Note that because this interpolation is performed over the NSS catalog, the interpolation is not defined for all pairs of ($N_{vis}$, $G$); the relationship would need to be extrapolated in order to estimate $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ at fainter magnitudes.
The main concern with this approach is that by using the NSS catalog itself to estimate the errors, this will propagate whatever selection effects we are trying to characterize into the prediction itself. However, since the final forward model provides reasonable agreement with observations, we conclude that this approach is acceptable, if not ideal.
4.3 Lack of sensitivity at $P\approx 1$ year and lack of solutions with $P>550$ days
Figure 5: Empirical function that removes orbital periods around 1 year and at long orbital periods $P_{orb}\gtrsim 550$ days. This empirical filter is applied after the $\texttt{ruwe}>1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters. The data used to make this figure are available through “Data behind the Figure”.
The black outline in the topmost 1-D histogram in Figure 7 shows the distribution of all NSS astrometric orbital periods in Gaia. There is a clear dearth of solutions around 1 year because the astrometric signature of annual parallax is degenerate with $P_{orb}=1$ year. The 1-D histogram also shows a steep drop-off in the number of solutions at longer orbital periods. This is expected for $P_{orb}\gtrsim 1000$ days due to Gaia DR3 only containing 34 months of data. However, the number of solutions rapidly drops around $P_{orb}\approx 550$ days, which is much shorter than the Gaia DR3 temporal baseline.
@@ 4.3 Lack of sensitivity at $P\approx 1$ year and lack of solutions with $P>550$ days
Forward modeling indicates that the ruwe  and parallax over parallax error filters do not cause the peak of the period distribution to be less than 2000 days (§5.1.2). We hypothesize that the drop-off in solutions at 550 days is primarily due to the astrometric processing cascade. As described in §2.3, acceleration solutions are tried before orbital solutions in the model cascade, and if an acceptable acceleration solution is found, it is kept and an orbital solution is not searched for, even if an orbital solution would have resulted in a better fit. Thus, long-period binaries are more likely to be satisfactorily fit with an acceleration model than binaries with orbital periods short enough to be covered multiple times over DR3.

### v2

@@ 4 Empirical modeling of Gaia NSS model cascade
With a binary population model in hand, the next step is to reproduce the modeling cascade described in §2. In this section, we describe how we estimate and implement the filter that selects astrometric NSS candidates ($\texttt{ruwe}>1.4$), the effect of the acceleration models in the cascade, and the filters that select the accepted orbital solutions ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ and $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{days}})$). Our general framework is to estimate the quantities of interest analytically, then use the simulations of El-Badry et al. (2024) to calibrate these estimates as a function of intrinsic orbital and binary parameters (e.g., period, ecccentricity, photocenter orbit size).
4.1 El-Badry et al. (2024) simulations
El-Badry et al. (2024) generate a synthetic population of Galactic binary stars, then forward model individual 1-D astrometric measurements of these sources according to Gaia’s scanning law and noise properties. They then model these 1-D observations according to the NSS astrometric cascade, fitting the simulated epoch data with the various NSS astrometric models and applying the various filtering criteria. El-Badry et al. (2024) find good agreement between the properties of their forward modeled catalog and the actual Gaia DR3 NSS astrometric orbits catalog.
Many of the underlying components of the population synthesis in El-Badry et al. (2024) are similar or identical to those described in this paper; they also use galaxia and COSMIC to simulate the population of binary stars within 2 kpc. Minor differences include the extinction ($A_{G}=2.3E(B-V)$ in this work vs. $A_{G}=2.8E(B-V)$ in El-Badry et al. 2024), number of binaries at birth within 2 kpc (170 million in this work vs. 195 million in El-Badry et al. 2024), and the white dwarf initial-final mass relation (Kalirai et al. 2008 in this work vs. Weidemann 2000 in El-Badry et al. 2024). However, we emphasize that the results we derive from El-Badry et al. (2024) in the rest of this section are independent of the details of their population synthesis. This is because we are considering the effects of the cascade relative to the initial catalog. As long as all relevant regions of parameter space are covered, then the results are independent of the specific assumptions used to generate the binary population.
4.2 ruwe  model
@@ 4.2.1 Analytic ruwe  prediction
$\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1}.$ (3)
@@
$\delta\theta=a_{0}\sqrt{1-\frac{\sin^{2}i}{2}-\frac{3+\sin^{2}i(\cos^{2}\omega-2)}{4}e^{2}},$ (4)
where $\omega$ is the argument of periapse. [^12] In Penoyre et al. (2020), $\omega$ is denoted $\phi_{\nu}$ and is called the azimuthal viewing angle of the binary. We have chosen to use the more standard notation of the Campbell elements. Note that the derivation of Equation 4 assumes that there are many observations sampling the full orbital period of the binary during the survey, such that the observations can be time-averaged over the orbit.
For an isotropic distribution of $\omega$, the average value of $\cos^{2}$$\omega$ is 0.5. Fixing $\cos^{2}$ $\omega$ to 0.5, the astrometric excess noise can be re-written
@@
Figure 4: Using the NSS astrometric orbits catalog, we validate the scaling relations between the approximation of the astrometric excess noise $\delta\theta\approx\sigma_{AL}\sqrt{\texttt{ruwe}^{2}-1}$ and the semimajor axis of the photocenter orbit $a_{0}$, eccentricity $e$, and orbital inclination $i$. The left column shows the approximation of $\delta\theta$ vs. $a_{0}$, with the color of the point denoting $|\cos i|$ (top panel) or eccentricity of the orbit (bottom panel). There is a linear trend between $a_{0}$ and $\delta\theta$ where the slope is less than unity, as expected. In addition, there is a clear gradient in the colors of the points: averaging over the orbit, face-on and circular orbits $(|\cos i|\approx 1,e\approx 0)$ have larger astrometric excess noise at fixed $a_{0}$ than edge-on and highly eccentric orbits $(|\cos i|\approx 0,e\approx 1)$. The right column is the same as the left column, but includes the dependence of $e$ and $i$ on $\delta\theta$, i.e. shows the approximation of $\delta\theta$ vs. $a_{0}\sqrt{(1+\cos^{2}i)(1-3e^{2}/4)}$. The trends in the colors of the points mostly disappear (the $|\cos i|$ and $e$ dependence), validating the scaling relation. There are still some minor trends visible in $|\cos i|$, but Gaia Collaboration et al. (2023b) showed that the sensitivity of circular orbits deviates slightly from the expected $1+\cos^{2}i$, so this is not surprising. Note that there are very few red points in the bottom panels because there are very few high-eccentricity orbits.
@@ 4.2.2 Empirical correction to analytic ruwe  prediction
The predicted astrometric excess noise in Equation 5 and ruwe  in Equation 6 will generally be overestimates of the astrometric excess noise and ruwe  reported by Gaia. The derivation of Equation 5 assumes that the binary signal has not been erroneously attributed to other astrometric terms (i.e., position, parallax, proper motion, or satellite pointing noise). Penoyre et al. (2020) predicted that this erroneous attribution would happen frequently, often lowering the reported astrometric excess noise and ruwe. With the NSS catalog, we confirm that the predicted astrometric excess noise $\delta\theta$ tends to be larger than the actual $\delta\theta$ in the catalog by $\approx 60-70\%$.
We calculate an empirical correction to rescale the analytic ruwe  prediction to match the ruwe  measured by fitting a single-star model to the simulated observations in the catalog of El-Badry et al. (2024). We use a catalog that includes a random subset of the birth population of binaries for which the orbital parameters are known and ruwe is calculated, but before the $\texttt{ruwe}>1.4$ filter has been applied. For each simulated source in this catalog, we calculate $\delta\theta_{analytic}$ (the expected astrometric excess noise using Equation 5) and $\delta\theta_{fit}$ (the astrometric excess noise inferred from the ruwe  measured in the fitting procedure; Equation 3). The ratio $\delta\theta_{fit}/\delta\theta_{analytic}$ is the rescaling factor required to correct the analytic prediction of $\delta\theta$ to account for the erroneous attribution that occurs in the fitting process. Within each bin, we only consider sources with $\sigma_{AL}<a_{0}$ (as otherwise their orbits are not detectable), and sources with $\texttt{ruwe}>1.01$ (as otherwise the predicted value of $\delta\theta$ will be imaginary).
In order to account for the dependence of this correction factor on the orbital period and number of observations, we split the simulated catalog solutions from El-Badry et al. (2024) into orbital period bins of 25 days, spanning 0 to 1500 days. We further subdivide each period bin based on the number of visibility periods used: $N_{vis}<25$, and $N_{vis}\geq 25$. These bins were chosen since $N_{vis}$ is a bimodal distribution for the NSS astrometric orbits catalog, with peaks at $N_{vis}=18$ and 27. For each $N_{vis}$ bin, a Gaussian is fit to the ratio $\delta\theta_{fit}/\delta\theta_{analytic}$. The mean of this Gaussian is taken to be the correction factor for that range of $P_{orb}$ and $N_{vis}$ (Figure 5). To determine the correction for a particular orbit, we interpolate across the $P_{orb}-N_{vis}$ grid to find the correction factor to $\delta\theta$, allowing the accurate prediction of ruwe  without actually having to generate and fit mock observations.
We treat viewer-dependent quantities (e.g., argument of periastron, inclination, time of passage through periastron) as nuisance parameters; our probabilistic function is effectively integrating over these quantities. Although there are selection effects due to viewing angle (e.g., face-on orbits are more likely to show a larger astrometric signal than an edge-on orbit), these are not intrinsic binary properties. We are not concerned with characterizing these detection biases since these are randomized across all binaries. This enables us to understand the trends analytically, and the qualitative and quantitative effects of the fitting and modeling procedures.
As shown in Figure 5, this correction is a strong function of orbital period. For $P_{orb}\gtrsim 1000$ days, the key assumption used to derive Equation 4 (having a fully-observed orbit) is violated; even for $P_{orb}\approx 1000$ days, this assumption may not hold if the orbit is not well-sampled for any reason (e.g., a highly eccentric orbit, or low number of visibility periods). At very long orbital periods $P_{orb}\gg 1000$ days, such a small fraction of the orbit will be covered by the Gaia DR3 temporal baseline that Equation 4 is no longer a valid approximation. The number of visibility periods used in the solution ($N_{vis}$) also affects the size of the erroneous attribution. For larger $N_{vis}$, the average size of the erroneous attribution is slightly smaller because more independent measurements better constrain the true orbit.
We emphasize that we only model orbital periods shorter than 1500 days. This is approximately the range over which there is a still a useful correlation between the prediction of Equation 4 and the true ruwe. This choice is justified because in the NSS catalog, only 1% of solutions have $P_{orb}>1500$ days; this is a negligible fraction of solutions compared to the other uncertainties in the forward modeling. We recommend using the methodology of El-Badry et al. (2024) and working through the full framework of epoch modeling when trying to understand $P_{orb}>1500$ day solutions.
Figure 5: Comparison of the astrometric excess noise $\delta\theta$ for sources in the synthetic catalog of El-Badry et al. (2024) as inferred by their ruwe, compared to the $\delta\theta$ as predicted by Equation 5. At longer orbital periods, the rescale factor gets smaller; when a orbit is only partially observed, it can be more easily be fit a single-star model, decreasing the astrometric excess noise. In addition, there are clear dips at orbital periods that are harmonics of 1 year or related to the scanning law. Finally, when the number of visibility periods is higher, the rescale factor is slightly larger (closer to unity) and the dips at the harmonics of 1 year are smaller. This is because when the orbit is better sampled, the solution is less prone to degeneracies that lead to incorrect solutions.
4.3 Removal of acceleration solutions candidates
@@
Although understanding the selection effects of the acceleration solutions catalog is not the focus of this work, it must be modeled because it precedes the orbital solution modeling in the cascade. This is because a source that has an acceleration solution accepted in the main processing is removed from being a candidate for an orbital solution (Figure 1). As described in §2.3, acceleration solutions are tried before orbital solutions in the model cascade, and if an acceptable acceleration solution is found, an orbital solution is not searched for, even if an orbital solution would have resulted in a better fit. Thus, long-period binaries are more likely to be satisfactorily fit with an acceleration model than binaries with orbital periods short enough to be covered multiple times over DR3, and are likely to not be included in the orbit catalog.
The intrinsic binary parameters that most strongly determine whether an acceleration solution is accepted are the orbital period $P_{orb}$ and the size of the photocenter orbit semimajor axis normalized by the measurement error $a_{0}/\sigma_{AL}$ [^13] Although a source’s positional uncertainty is determined by both its brightness and number of visibility periods, it is primarily a function of the brightness; thus, at the moment, we do not model the number of visibility periods.. Eccentricity does not significantly affect the acceptance probability; see Appendix C for details.
We use the simulated catalog of El-Badry et al. (2024) to construct an empirical probabilistic function to determine which sources will be candidates for orbital solutions. We start with all the NSS sources that pass the initial filtering step $(\texttt{ruwe}>1.4)$ and bin them by $P_{orb}$ and $a_{0}/\sigma_{AL}$. Within each $P_{orb}-a_{0}/\sigma_{AL}$ bin, we calculate the fraction of sources which are candidates for orbital solutions (sources where neither a 9-parameter or 7-parameter acceleration solution is accepted in the main processing). This is shown in Figure 6. To find the probability that a source with $\texttt{ruwe}>1.4$ is a candidate for an orbital solution, we interpolate across this $P_{orb}-a_{0}/\sigma_{AL}$ grid.
The shapes of the high and low probability regions in Figure 6 can be understood as follows. A low probability of trying an orbital model is equivalent to a high probability of accepting an acceleration solution, and vice versa. Acceleration solutions are unlikely to be accepted at short orbital periods $P_{orb}\lesssim 500$ days because Gaia DR3 will have covered at least two full orbits, and an acceleration model will generally be a very poor fit; thus, the probability of trying an orbital solution is very high for $P_{orb}\lesssim 500$ days.
At longer orbital periods $P_{orb}\gtrsim 500$ days, the probability of accepting an acceleration solution is maximal around the “sweet spot” of $\log_{10}(a_{0}/\sigma_{AL})\approx 0.8$ and $P_{orb}\approx 1300$ days. If $\log_{10}(a_{0}/\sigma_{AL})\lesssim 0.8$, then the detected acceleration will be noisy and less significant, making it less likely to pass the acceleration acceptance criterion on model significance. On the other hand, if $\log_{10}(a_{0}/\sigma_{AL})\gtrsim 0.8$, the astrometric signal will be much stronger, but the suitability of the acceleration model will depend on the orbital period. For fractionally covered orbits $P_{orb}\gtrsim 1000$ days, an acceleration model will be a good fit. However, for orbits $P_{orb}\approx 500-1000$ days where a full orbit or more has been covered, measurement noise is needed to mask the model mismatch. This explains the diagonal boundary separating the high and low probability regions for $P_{orb}\gtrsim 500$ days and $\log_{10}(a_{0}/\sigma_{AL})\gtrsim 0.8$ in Figure 6.
Figure 6: Probability than an orbit model is tried for an NSS candidate, as a function of orbital period $P_{orb}$ and the photocenter orbit size $a_{0}$, normalized by the scan error $\sigma_{AL}$. Nearly all sources with orbital periods $P_{orb}\lesssim 500$ days are candidates for orbital solutions. However, for $P_{orb}\gtrsim 500$ days, the longer orbit solutions are increasingly unlikely to be candidates for an orbital solution, especially those with less significant photocenter orbit sizes. This is because these solutions are lost to the acceleration portion of the model cascade.